Lydia has money in two savings accounts. One rate is 6% and the other is 14%. If she has $950 more in the 14% account and the total interest is $318, how much is invested in each savings account?
Answer by mananth(16949)

let amount invested in 6% account be x
in 14% account it will be x+950

0.06x+0.14(x+950)= 318
0.06x + 0.14x + 133=318
0.2x=185
x= 185/0.2
= $925 invested in 6% account
925+950= 1875.
$ 1875 is invested in 14% account